OpenStudy (jthec):

6. The oxidation of malate by NAD+ to form oxaloacetate is a highly endergonic reaction under standard conditions (G’o = +7.0 kcal/mol). The reaction proceeds readily under physiological conditions: (b) Assuming an [NAD+]/[NADH] ratio of 8 and a pH of 7, what is the lowest [malate]/[oxaloacetate] ratio at which oxaloacetate can be formed from malate?

OpenStudy (anonymous):

whats the question??

OpenStudy (jthec):

what is the lowest [malate]/[oxaloacetate] ratio at which oxaloacetate can be formed from malate?

OpenStudy (caozeyuan):

Ok so this is equilibrium question

OpenStudy (caozeyuan):

\[K=\frac{ \left[ oxaloacetate \right] \left[ NADH\right] }{ \left[ malate \right] \left[ NAD ^{+}\right] }\]

OpenStudy (caozeyuan):

you want K>=1 so that action proceed to the right

OpenStudy (jthec):

yeah

OpenStudy (caozeyuan):

[NAD+]/[NADH]=8 so for K to be 1, [malate]/[oxaloacetate] should be 1/8